OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2)

Nnesha (nnesha):

first write this equation 3x-y=-3 in slope intercept form :)

Nnesha (nnesha):

do you know who to solve for y ?? 3x -y =-3 slope intercept form is y = mx + b :)

OpenStudy (anonymous):

Would the slope intercept form be y=3x-3?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

well slope is right but y-intercept suppose to be positive bec y is negative at left side so you hve to divide